OpenStudy (anonymous):

wanna make sure I am write on this question Rewrite with only sin x and cos x. sin 3x I got 2 cos2x sin x + sin x - 2 sin3x

OpenStudy (anonymous):

\[\sin 3x = 3\sin x - 4\sin^3x\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

this is all my choices

OpenStudy (anonymous):

2 sin x cos2x + cos x 2 sin x cos2x + sin3x sin x cos2x - sin3x + cos3x 2 cos2x sin x + sin x - 2 sin3x

OpenStudy (anonymous):

sin 3x = sin (2x+x) = sin 2x cos x+ cos2x sinx = 2sinx cos x cos x+ cos2x sinx \[= 2sinx \cos ^2x+ (1-\sin^2x) sinx\] \[= 2sos ^2x inx c+ sinx-\sin^2x sinx \] \[= 2\cos ^2x sinx + sinx-\sin^3x \]

OpenStudy (anonymous):

@llkramer

OpenStudy (anonymous):

so I was right?

OpenStudy (anonymous):

@llkramer how have you got 2 sin3x?

OpenStudy (anonymous):

oh so I was wrong then its this one sin x cos2x - sin3x + cos3x I knew it was this one or the other one because the other choices the xs arnt correct right

OpenStudy (anonymous):

or no it was the coeff that weren't right

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Ok let us solve it again.

OpenStudy (anonymous):

sin 3x = sin (2x+x) = sin 2x cos x+ cos2x sinx Let us know the reasons first.

OpenStudy (anonymous):

since sin2x = 2 sinx cosx and \[\cos2x = 1- 2\sin^2x\] Now substituting these values, we find.

OpenStudy (anonymous):

sin 3x = sin (2x+x) = sin 2x cos x+ cos2x sinx \[= 2 sinx \times cosx \times \cos x+ (1-2\sin^2x) sinx\] \[= 2 sinx \times \cos^2x + sinx-2\sin^2x \times sinx \] \[= 2 sinx \times \cos^2x + sinx-2\sin^3x \] \[= 2 \cos^2x sinx + sinx-2\sin^3x \]

OpenStudy (anonymous):

I was littele bit wrong and thats why was not able to get the answer.

OpenStudy (anonymous):

@llkramer

OpenStudy (anonymous):

haha so I was write in the beginning xD?

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

@llkramer

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

so I was write in the begging?

OpenStudy (anonymous):

i'm also here

OpenStudy (anonymous):

so I was write

OpenStudy (anonymous):

Yes, correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sweet thank you

OpenStudy (anonymous):

actually i used wrong formula in between

OpenStudy (anonymous):

just submitted it and I got it right

OpenStudy (anonymous):

Congrats